# The Devil behind the mask: An emergent safety vulnerability of Diffusion LLMs

### v1

@@ A.1 Victim Models
  • LLaDA-1.5101010https://huggingface.co/GSAI-ML/LLaDA-1.5 (Zhu et al., 2025) introduces Variance-Reduced Preference Optimization (VRPO) to address the key bottleneck in the alignment of diffusion-based language models: unstable gradient updates caused by high-variance ELBO approximations. The method employs optimized Monte Carlo sampling that prioritizes diverse timesteps over mask patterns, along with antithetic noise coupling for preference comparisons to cancel estimation errors, and increased sample sizes for greater stability. VRPO enables robust reward-model fine-tuning that outperforms SFT-only baselines on reasoning tasks, demonstrating that RLHF-style alignment is viable for diffusion models through effective variance control.
  • Dream-v0-Instruct-7B [^11] https://huggingface.co/Dream-org/Dream-v0-Instruct-7B (Ye et al., 2025) specializes in reasoning tasks through diffusion modeling. Unlike standard approaches, it strategically initializes weights from autoregressive models and introduces adaptive per-token noise scheduling, dynamically adjusting difficulty levels based on contextual clues. These innovations enable it to rival larger AR models like LLaMA3-8B while maintaining training efficiency, demonstrating how diffusion models can effectively learn complex reasoning patterns.
  • MMaDA-8B-MixCoT [^12] https://huggingface.co/Gen-Verse/MMaDA-8B-MixCoT (Yang et al., 2025) features a modality-agnostic diffusion architecture and a unified probabilistic formulation, eliminating modality-specific components. A mixed long CoT fine-tuning strategy enhances instruction-following and stabilizes CoT generation over MMaDA-8B-Base [^13] https://huggingface.co/Gen-Verse/MMaDA-8B-Base.
A.2 Benchmarks
@@
  • JailbreakBench (Chao et al., 2024a) is an open-source benchmark for evaluating jailbreak attacks on large language models, addressing key challenges in standardization and reproducibility. It features (i) a continuously updated repository of adversarial prompts, (ii) a curated dataset of 100 policy-violating behaviors, (iii) a standardized evaluation framework with defined threat models and scoring metrics.
@@ A.3 Attack Baselines
A.4 Defenses
In our study, we incorporate two defense methods, whose implementation details are provided in Tables 4 and 5, respectively.
@@ A.5 Our Prompt Template
Appendix B Evaluation Metrics
@@ B.2 StrongREJECT Score
B.3 Harmfulness Score.
We also adopt the GPT-judged Harmful Score (HS) as our primary evaluation metric. Specifically, GPT-4o is employed as the evaluator to assess the victim model’s responses to adversarial prompts based on both harmfulness and relevance. The harmful score ranges from 1 to 5, where a score of 1 indicates that the model either refused to respond or produced a response that is harmless or irrelevant, while a score of 5 denotes a highly harmful or relevant response. In our experiments, responses assigned HS=5 are regarded as successful jailbreaks. As shown in Tables 8 and 9, we follow prior work and adopt the same GPT-judge prompt for consistency.
@@
Appendix C More Experimental Results
In our extended experiments, we comprehensively evaluated the defensive capabilities of dLLMs and autoregressive LLMs against various jailbreak attacks using three benchmarks: HarmBench, JailBreakBench, and StrongREJECT. As shown in Figure 8, dLLMs exhibit an evaluator-based ASR (ASR-e) that is comparable to or even lower than that of autoregressive LLMs under the AIM and PAIR attacks. This trend is consistent with the findings presented in Figure 3 in the main text. Meanwhile, the results in Figure 9 indicate that dLLMs generally exhibit comparable or slightly better initial resistance to zero-shot attacks compared to autoregressive LLMs, as evidenced by lower keyword-based and evaluator-based attack success rates (ASR-k and ASR-e). In summary, our comprehensive evaluation across multiple benchmarks and attack scenarios reveals that dLLMs often match or surpass those of autoregressive LLMs in resisting existing jailbreak attack methods.
Appendix D Limitations and Future Works
While our study uncovers critical vulnerabilities in diffusion-based large language models (dLLMs), several limitations remain to be addressed. First, the absence of white-box attack baselines, such as GCG (Zou et al., 2023), limits the depth of our analysis. Future work should incorporate these approaches to enable a more fine-grained examination of internal vulnerabilities. Second, our experiments focus on a limited set of models; extending the evaluation to additional dLLMs, including DiffuCoder (Gong et al., 2025) and LongLLaDA (Liu et al., 2025), would enhance the generality of our conclusions and further validate DiJA’s effectiveness. In future works, there is a pressing need to develop safety alignment techniques specifically designed for the unique generative processes of dLLMs. Meanwhile, exploring the safety of multi-modal (Li et al., 2025; You et al., 2025) and unified dLLMs (Yang et al., 2025), particularly in tasks involving image and video generation (Chen et al., 2024a) and multi-modal understanding, is essential for a more comprehensive understanding of their security implications.

### v2

@@ C.1 Victim Models
  • LLaDA-1.5121212https://huggingface.co/GSAI-ML/LLaDA-1.5 (Zhu et al., 2025) introduces VRPO, a variance-reduced optimization that stabilizes diffusion model alignment and enables effective RLHF-style fine-tuning, outperforming SFT-only baselines.
  • Dream-v0-Instruct-7B [^13] https://huggingface.co/Dream-org/Dream-v0-Instruct-7B (Ye et al., 2025) is a diffusion-based model focused on reasoning tasks. It initializes from autoregressive weights and uses adaptive noise scheduling, allowing it to match larger AR models like LLaMA3-8B in performance while remaining efficient.
  • MMaDA-8B-MixCoT [^14] https://huggingface.co/Gen-Verse/MMaDA-8B-MixCoT (Yang et al., 2025a) features a modality-agnostic diffusion architecture and a unified probabilistic formulation, eliminating modality-specific components. A mixed long CoT fine-tuning strategy enhances instruction-following and stabilizes CoT generation over MMaDA-8B-Base [^15] https://huggingface.co/Gen-Verse/MMaDA-8B-Base.
  • DiffuCoder-7B-Instruct [^16] https://huggingface.co/apple/DiffuCoder-7B-Instruct (Gong et al., 2025) is a 7B discrete diffusion language model for code, trained on ~130B code tokens and instruction-tuned for coding tasks, featuring any-order generation via global sequence denoising rather than left-to-right decoding.
  • DiffuCoder-7B-cpGRPO [^17] https://huggingface.co/apple/DiffuCoder-7B-cpGRPO (Gong et al., 2025) is the RL fine-tuned variant of DiffuCoder-7B using coupled-GRPO, a diffusion-native reinforcement learning scheme with coupled sampling that boosts code-generation performance; it retains the 7B discrete diffusion architecture with any-order generation via global sequence denoising.
  • Dream-Coder-v0-Instruct-7B [^18] https://huggingface.co/Dream-org/Dream-Coder-v0-Instruct-7B (Xie et al., 2025) is a 7B discrete diffusion language model for code with emergent any-order generation; it adapts a pretrained autoregressive checkpoint to a diffusion objective (continuous-time weighted cross-entropy) and is instruction-tuned with additional RL using verifiable rewards.
C.2 Benchmarks
@@
  • JailbreakBench (Chao et al., 2024a) is an open-source benchmark for evaluating jailbreak attacks on large language models, addressing key challenges in standardization and reproducibility. It features (i) a continuously updated repository of adversarial prompts, (ii) a curated dataset of 100 policy-violating behaviors, and (iii) a standardized evaluation framework with defined threat models and scoring metrics.
@@ C.3 Attack Baselines
  • GCG (Zou et al., 2023) is a gradient-based white-box jailbreak attack that iteratively edits prompt tokens to maximize an adversarial objective. It uses gradient information to greedily update coordinates (tokens), refining the adversarial prompt over iterations.
C.4 Defenses
In our study, we incorporate two defense methods, whose implementation details are provided in Tables 13 and 14, respectively.
@@ C.5 Our Prompt Template
C.6 Model Interfaces and Mask Syntax
To clarify the portability of DiJA and the assumed I/O interfaces, we provide a detailed mapping of the mask token syntax for the evaluated dLLMs and other representative models in Table 16. Currently, most state-of-the-art dLLMs are open-source, allowing users direct access via frameworks such as HuggingFace Transformers. In this setting, inputs are not sanitized by a serving stack, enabling the direct injection of raw mask tokens. For closed-source models (e.g., Mercy Diffusion, Seed Diffusion), raw mask injection is currently unavailable to the public. However, since remasking and re-generation are core competitive advantages of dLLMs over autoregressive models (as shown in Figure 1), disabling mask inputs would strip these models of their native capabilities. Therefore, we anticipate that even restricted interfaces will likely expose mechanisms (e.g., specific API endpoints for infilling or editing) that correspond to the mask placements exploited by DiJA.
Table 16: Mapping of mask token syntax and access interfaces for representative dLLMs. ✓ denotes open-source availability, while $\times$ denotes closed-source.
Model Family Open Source Mask Token Access Interface / Checkpoint
LLaDA-8B Family <|mdm_mask|> https://huggingface.co/GSAI-ML/LLaDA-8B-Base
LLaDA-MoE <|mask|> https://huggingface.co/inclusionAI/LLaDA-MoE-7B-A1B-Base
LLaDA-2.0 <|mask|> https://huggingface.co/inclusionAI/LLaDA2.0-mini-preview
Dream-7B Family <|mask|> https://huggingface.co/Dream-org/Dream-v0-Base-7B
MMaDA Family <|mask|> https://huggingface.co/Gen-Verse/MMaDA-8B-MixCoT
DiffuCoder Family <|mask|> https://huggingface.co/apple/DiffuCoder-7B-Instruct
SDAR Family <|MASK|> https://huggingface.co/JetLM/SDAR-8B-Chat
Mercy Diffusion $\times$ - https://www.inceptionlabs.ai
Seed Diffusion $\times$ - https://studio.seed.ai/exp/seed_diffusion
Appendix D Evaluation Metrics
@@ D.2 StrongREJECT Score
D.3 Harmfulness Score.
We also adopt the GPT-judged Harmful Score (HS) as our primary evaluation metric. Specifically, GPT-4o is employed as the evaluator to assess the victim model’s responses to adversarial prompts based on both harmfulness and relevance. The harmful score ranges from 1 to 5, where a score of 1 indicates that the model either refused to respond or produced a response that is harmless or irrelevant, while a score of 5 denotes a highly harmful or relevant response. In our experiments, responses assigned HS=5 are regarded as successful jailbreaks. As shown in Tables 19 and 20, we follow prior work and adopt the same GPT-judge prompt for consistency.
D.4 Human Validation of Harmfulness Score
To bolster the soundness of the reported Harmfulness Score (HS) and address potential concerns regarding evaluator dependence, we conducted a human validation experiment to cross-check the GPT-4o judge with human evaluators. We randomly sampled a subset of 100 responses (25 from each of the four victim dLLMs) generated under the DiJA attack. Two independent human evaluators scored each response on a scale of 1 to 5, strictly following the same Harmfulness Score rubric used by GPT-4o (as detailed in Table 19). We then validated the alignment between Human and GPT-4o judgments on two levels:
  • Harmfulness Detection (Binary Agreement): To verify if GPT-4o correctly identifies harmful content, we binarized the scores with a threshold of 3. Scores $\geq 3$ were classified as “Harmful”, and scores <3<3 as “Safe/Refusal”. We calculated the Agreement Rate and Cohen’s Kappa ($\kappa$) on this binary classification.
  • Score Consistency (Pearson Correlation): We calculated the Pearson correlation coefficient ($r$) between the raw 1-5 scores to ensure GPT-4o accurately reflects the granularity of harmfulness across the severity spectrum.
Table 18: Agreement between human evaluators and GPT-4o judge on harmfulness assessment (N=100N{=}100). High agreement rates and correlation coefficients indicate GPT-4o is a reliable proxy for human evaluation.
Pair Binary Agreement Cohen’s Kappa ($\kappa$) Pearson Correlation ($r$)
Human A vs. GPT-4o 94.0% 0.88 0.92
Human B vs. GPT-4o 93.0% 0.86 0.90
Human A vs. Human B 96.0% 0.92 0.95
The results, summarized in Table 18, demonstrate high consistency. The Binary Agreement exceeds 93% with a Cohen’s Kappa >0.8>0.8, indicating strong agreement in distinguishing harmful from safe responses. Furthermore, the Pearson Correlation >0.9>0.9 confirms that GPT-4o’s scoring trend is highly synchronized with human intuition. These findings validate the reliability of the Harmfulness Score (HS) reported in our main experiments.
@@ D.4 Human Validation of Harmfulness Score
Appendix E Limitations and Future Works
While our study uncovers critical vulnerabilities in diffusion-based large language models (dLLMs), several limitations remain to be addressed. First, exploring the safety of multi-modal (Li et al., 2025; You et al., 2025) and unified dLLMs (Yang et al., 2025a), particularly in tasks involving image and video generation (Chen et al., 2024a) and multi-modal understanding, is essential for a more comprehensive understanding of their security implications. Meanwhile, although we explored an architecture-aware alignment method for dLLMs (Appendix A), significant opportunities for future research remain. Key directions include: (i) evaluating the generalization and robustness of the proposed refusal-aware alignment against unseen or modified jailbreak patterns; (ii) quantifying any potential degradation in utility on more benign tasks, particularly those involving interleaved mask-text prompts (as shown in Figure 1). (iii) Privacy risks (Yang et al., 2024b; Wen et al., 2024) and watermarking defense issues (Gloaguen et al., 2025) in dLLMs are also important and worthwhile directions to explore.
Appendix F Use of LLMs
In this study, we only utilized large language models (LLMs) to perform grammar checking and to polish certain sentences for improved clarity and fluency, without altering the original meaning of the text. The core work, including idea conception, experimental design, and data analysis, was conducted by the authors.